OpenStudy (anonymous):

Solve the equation. x+1/6=6 A.) x= -6 1/6 B.) x= -5 5/6 C.) x= 5 5/6 D.) x= 6 1/6

OpenStudy (anonymous):

Subtract 1/6 to both sides: \(6 - \dfrac{1}{6} \rightarrow \dfrac{36}{6} - \dfrac{1}{6} \rightarrow \dfrac{36-1}{6}\) Can you subtract 36 - 1? @Username_00